show that y=c1e^x+c2e^2x is the general solution of y′′−3y′+2y=0 on any interval and find the particular solution for which y(0)=-1 and y'(0)=1


1
Expert's answer
2022-01-31T15:04:34-0500

"y''\u22123y'+2y=0"

"D\u00b2y-3Dy+2y=0"

"(D\u00b2-3D+2)y=0"


This is a homogeneous equation. Thus, we have that

"m\u00b2-3m+2 = 0"

"(m-2)(m-1) = 0"

"=>m=2, 1"


The general solution is


"y=C_{1} e^x+C_{2}e^{2x}"


"y'=C_{1} e^x+2C_{2} e^{2x}"


But "y(0)=-1" and "y'(0)=1"

"=>"

"C_1+C_2=-1"

"C_1+2C_2=1"

Solving both equations simultaneously, we have that

"\u2234C_1 = -3" and "C_2 = 2"


Hence, the particular solution is

"y=2e^{2x}-3e^x"